OpenStudy (anonymous):

Find the exact value of tan 2x if tan x = 7/11and PI is less than or equal to x which is less than or equal to 3pi/2

OpenStudy (anonymous):

have you come across compound angle formulae yet?

OpenStudy (anonymous):

I was probably supposed to... but no.

OpenStudy (anonymous):

well there's a crude but quick way of doing this, or the compound angle way which is probably the official way

OpenStudy (anonymous):

does this look familiar? : \[\tan2x = \frac{2tanx}{1-2\tan^2{x}}\]

OpenStudy (anonymous):

Yws.

OpenStudy (anonymous):

ok, well substitute into that formula tanx = 7/11

OpenStudy (anonymous):

tan2x = 2(7/11)/(1-2(49/121))

OpenStudy (anonymous):

= (14/11)/(23/121) =(14*121)/(23*11)

OpenStudy (anonymous):

etc etc

OpenStudy (anonymous):

I got it..

OpenStudy (anonymous):

:D good good